# Value function interference and greedy action selection in value-based multi-objective reinforcement learning

### 09 Feb 2024 v1

@@ 4 Conclusion
Value function interference is an issue affecting value-based approaches to multi-objective reinforcement learning, which has not previously been identified in the MORL literature. These methods usually learn Q-values which are estimates of the expected mean future vector return, and in some cases this is not sufficient information to establish the optimal policy. This is particularly evident when trying to learn ESR-optimal policies for stochastic environments, but this paper has also demonstrated that it is possible for value function interference to arise in the context of SER-optimal policies for deterministic environments. We have shown empirically that in the latter case, modifying the core Q-learning algorithm to use non-random tie-breaking when two actions have equivalent utility can ameliorate, but not totally eliminate, the impact of value function interference. More generally, we advocate the use of alternative approaches such as distributional forms of value-based MORL or policy-search MORL methods to eliminate or avoid this issue.

### 22 Apr 2026 v2

@@ 4 Conclusion
This paper has identified two previously unreported issues which may impact on the performance of value-based MORL algorithms, and provided empirical evidence of this impact on tabular multi-objective Q-learning when applied in combination with a non-linear utility function:
  • Value function interference is an issue arising from the use of vectors to represent the expected future multi-objective return. In some cases this is not sufficient information to establish the optimal policy. This is particularly evident when trying to learn ESR-optimal policies for stochastic environments, but this paper has also demonstrated that it is possible for value function interference to arise in the context of SER-optimal policies for deterministic environments. We have shown empirically that in the latter case, modifying the core Q-learning algorithm to use non-random tie-breaking when two actions have equivalent utility can ameliorate, but not totally eliminate, the impact of value function interference.
  • It is well established that value-based RL algorithms can be prone to overestimation bias when used in conjunction with function approximation. While this problem is not unique to MORL, the results in Section 3 highlight that MORL methods may be particularly sensitive to overestimated Q-values when applied with non-linear utility functions.
We suggest that one promising course of action to address value function interference is to incorporate concepts from distributional RL into MORL, as pioneered by Hayes et al. (2021),Reymond et al. (2023) and Röpke et al. (2023). If the agent learns a distribution over the vector returns for each state-action, then this can be combined with the utility function so as to estimate the ESR.
For example for the MOMAB in Figure 2 assume the agent has correctly learned that the distribution of future returns for action $a_{1}$ is (7,-1, -5) with $p=0.5$ and (7,-5, -1) with $p=0.5$. From this information the agent can correctly estimate the ESR utility of $a_{1}$ as the probabilistically-weighted sum of those return vectors.
$\begin{split}U_{ESR}(a_{1})=0.5U(7,-1,-5)+0.5U(7,-5,-1)\\ =0.5*9+0.5*9\\ =9\end{split}$ (10)
Another possible solution is to adopt the approach of Cai et al. (2023). This involves a form of potential-based reward shaping derived by using the multi-objective utility function to scalarise the accumulated vector return on each time-step, and using a conventional RL algorithm to learn a policy from those scalar rewards. This means the agent is directly learning the ESR value, and so eliminates the possibility of value-function interference. As it is learning a scalar return, this approach may also mitigate against the sensitivity of non-linear utilities to overestimation in Q-values. However it no longer supports some of the benefits of MORL, such as multi-policy learning, or the increased explainability offered by vector Q-values (Zhan and Cao, 2019; Dazeley et al., 2023).